# ZipVoice: Fast and High-Quality Zero-Shot Text-to-Speech with Flow Matching

### v2

@@ Abstract
Existing large-scale zero-shot text-to-speech (TTS) models deliver high speech quality but suffer from slow inference speeds due to massive parameters. To address this issue, this paper introduces ZipVoice, a high-quality flow-matching-based zero-shot TTS model with a compact model size and fast inference speed. Key designs include: 1) a Zipformer-based flow-matching decoder to maintain adequate modeling capabilities under constrained size; 2) Average upsampling-based initial speech-text alignment and Zipformer-based text encoder to improve speech intelligibility; 3) A flow distillation method to reduce sampling steps and eliminate the inference overhead associated with classifier-free guidance. Experiments on 100k hours multilingual datasets show that ZipVoice matches state-of-the-art models in speech quality, while being 3 times smaller and up to 30 times faster than a DiT-based flow-matching baseline. Codes, model checkpoints and demo samples are publicly available. [^1] https://github.com/k2-fsa/ZipVoice
@@ I Introduction
Recently, flow-matching [12], a variant of diffusion models, has been widely adopted in TTS [2, 3, 4, 13, 14] due to its reduced sampling steps compared to previous diffusion-based models. A representative flow-matching-based zero-shot TTS system is E2-TTS [3]. E2-TTS addresses text-speech alignment by padding text tokens with filler tokens to match the speech length, and letting the Transformer [15]-based flow-matching decoder to learn the alignment implicitly. This approach significantly simplifies the system architecture and training process by obviating the need for token-level duration prediction. Nevertheless, achieving satisfactory performance with E2-TTS still necessitates a large model size and dozens of sampling steps, which inevitably results in slow inference.
To address the aforementioned challenges, we introduce ZipVoice, a high-quality zero-shot TTS model with a compact model size and fast inference speed. Built upon the flow-matching framework for its simplicity and superior performance, ZipVoice ensure efficient and high-quality speech generation with following designs. Firstly, Zipformer [16] is incorporated as the backbone of flow-matching decoder to ensure adequate model capacity within a limited parameter budget. We demonstrate that the Zipformer architecture, although originally designed for automatic speech recognition (ASR), is also well-suited for TTS tasks. Secondly, since the practice of padding text tokens with filler tokens in E2-TTS results in suboptimal speech-text alignment and compromised intelligibility [4], to ensure the intelligibility of ZipVoice, on the one hand, we propose an average upsampling approach that assumes uniform token durations within a sentence. This straightforward strategy preserves system simplicity by eliminating the need for token-level duration prediction, while maintaining stable alignment between speech and text. On the other hand, a Zipformer-based text encoder is employed to extract better text token representations. Third, flow-matching-based TTS models typically require many sampling steps to generate high-quality speech, and the widely used classifier-free guidance (CFG) [17] strategy further introduces an additional inference pass without condition in each sampling step. To mitigate this, we design a flow distillation method to distill the pre-trained flow-matching model, thereby reducing the number of sampling steps and avoiding the extra computational overhead introduced by CFG.
Figure 1: Illustration of the ZipVoice model architecture.
@@ II-A Preliminary: Flow-Matching
After training with the CFM loss, we can generate samples by solving the ordinary differential equation (ODE). Using the Euler solver as an example, starting from an initial sample $x_{0}\sim p_{0}$ (e.g., Gaussian noise), we can iteratively updates the sample towards the target distribution $p_{1}$. For a discrete time sequence $10=t_{0}<\dots<t_{k}<\dots<t_{K}=$ 1, the Euler solver updates the sample at each step $k$ as:
@@ II-B Overview
The architecture of the ZipVoice model is illustrated in Fig. 1. It consists of two components: a text encoder and a flow-matching decoder, both employing Zipformer as the backbone. We explain the advantage of this design in subsection II-C.
@@
To enable zero-shot TTS capabilities, we adopt the speech infilling task following [2]. Specifically, a binary temporal mask $m\in\{0,1\}^{D\times T}$, where 1 denotes masked positions, is applied to the speech feature $x_{1}\in\mathbb{R}^{D\times T}$. The model is trained to reconstruct $m\odot x_{1}$ given the speech condition $(1-m)\odot x_{1}$, the interpolated noisy speech feature $x_{t}=(1-t)x_{0}+tx_{1}$, and the text condition $z$. As shown in Fig. 1, these three inputs have the same length and are concatenated along the feature dimension to form the input to the flow-matching decoder.
@@ II-C ZipVoice architecture with Zipformer Backbone
ZipVoice comprises a Zipformer-based text encoder and a Zipformer-based flow-matching decoder, with the latter housing the majority of the model’s parameters. Zipformer is an encoder structure designed for automatic speech recognition (ASR). It distinguishes itself from the standard Transformer structures through three key designs: U-Net-like architecture, convolution modules, and attention weight reuse mechanism.
Zipformer is well-suited as the flow-matching decoder backbone for the following reasons. Firstly, the U-Net architecture, which processes feature representations at varying resolutions, is widely recognized as an effective inductive bias in diffusion models [18, 19]. Secondly, convolutional neural networks (CNNs) are proficient in capturing fine-grained local feature patterns, complementing the Transformer’s strength in modeling long-range global dependencies. In TTS tasks, the nature of speech leads to strong correlations between adjacent hidden states [20, 21], making CNNs particularly well-suited for enhancing modeling capacity. Finally, Zipformer improves parameter efficiency by reusing attention weights across three modules within a single layer: two self-attention modules and a non-linear attention (NLA) module. This design not only reduces parameter count by sharing projection layers of query and key, but also enhances computational efficiency via attention weight sharing. Therefore, while originally engineered for ASR, these attributes endow Zipformer with ideal compatibility as the backbone of flow-matching decoders.
Although some recent flow-matching models [3, 4] omit dedicated text encoders by feeding text embeddings directly into the flow-matching decoder, we find that a well-designed text encoder can improve speech intelligibility. Our Zipformer-based text encoder retains convolutional modules and attention weight reuse while omitting the U-Net structure, aligning with hybrid CNN-Transformer designs in prior works [22, 23].
@@ II-D Speech-Text Alignment with Average Upsampling
Despite its simplicity, this approach suffers from inaccurate alignment [4] and slow convergence  [3]. F5-TTS addresses this by incorporating a ConvNeXt [26] module to refine the padded text condition. ZipVoice instead applies a parameter-free average upsampling strategy to address this issue, under the intuitive assumption of uniform token durations, i.e., each token has identical duration within a sentence. Specifically, when there are $N$ text tokens and $T$ frames speech features frames, the duration of each text token is computed with:
@@
[…]
Each text embedding is repeated $d$ times and the length of the text feature is expanded from $N$ to $d\cdot N$. If $T>d\cdot N$, the expanded text features is further padded with $T-d\cdot N$ filler embeddings. The final text feature $z\in\mathbb{R}^{F\times T}$ is used as the text condition of the flow-matching decoder.
@@ II-E Speedup ZipVoice with Flow Distillation
Given a pre-trained TTS model $\theta^{T}$ as the teacher, we initialize the student model $\theta^{S}$ with the parameters of $\theta^{T}$. Since we want the student model to benefit from the CFG inference while avoiding the additional model evaluation of CFG, we modify the student model to be conditioned on the CFG strength [27]. Specifically, the CFG strength $\omega$ is first processed through a Fourier embedding and a linear layer, and then integrated into the model, which is similar to how the time-step is incorporated into the flow-matching decoder.
@@
The two step sizes, $t_{\text{mid}}-t$ and $t_{\text{dest}}-t_{\text{mid}}$ are uniformly sampled from $[0,\Delta t_{\text{max}}]$. $\omega$ is also uniformly sampled between a given interval $[\omega_{\text{min}},\omega_{\text{max}}]$. We empirically find that dynamically sampling these values yields better performance than using fixed values.
Then, the teacher vector felid is computed as:
@@ II-F Inference Strategy
We design a time-dependent CFG strategy to achieve a better trade-off between speaker similarity and intelligibility: in early NFEs, only the text condition is dropped for unconditional predictions, while in later steps, both text and audio conditions are dropped. This strategy aims to focus more on content-related features in the early steps.
This strategy is exclusively applied to the non-distill version, ZipVoice. For ZipVoice-Distill, the teacher vector field constructed with ZipVoice has incorporated this strategy during flow distillation. Thus, ZipVoice-Distill inherits this property without requiring explicit inference-time adjustments.
TABLE I: Evaluation results of zero-shot TTS models trained on large-scale datasets. The boldface denotes the best result. * Results from papers. §$\S$ Unofficial implementations. $\dagger$ Inference with official checkpoints. $\ddagger$ trained by us with official codes.
@@ IV-B Model
ZipVoice consists of a Zipformer-based text encoder and a Zipformer-based flow-matching decoder. The text encoder consists of 4 Zipformer layers, each having an encoder dimension of 192, a feedforward dimension of 512. For the flow-matching decoder, it contains 5 stacks of Zipformer encoder layers operating at [1×,2×,4×,2×,1×][1$\times$,2$\times$,4$\times$,2$\times$,1$\times$][ 1 × , 2 × , 4 × , 2 × , 1 × ] downsampling rates and with $[2,2,4,4,4]$ layers, respectively. Each layer in the flow-matching decoder has a decoder dimension of 512 and a feedforward dimension of 1536. The total number of parameters of ZipVoice models are around 123M.
@@ V Experimental Results
V-A Comparison with Baseline models on large-scale datasets.
@@
V-B Comparison with Baseline models on small-scale datasets.
TABLE III: Evaluation results of zero-shot TTS models trained on small-scale datasets. The boldface denotes the best result. §$\S$ Unofficial implementations. $\dagger$ Inference with official checkpoints. $\ddagger$ trained by us with official codes.
@@ V-C Ablation experiments on the ZipVoice structure
TABLE IV: Ablation studies on the ZipVoice architecture
@@ V-D Ablation experiments on the Zipformer backbone
We conduct detailed ablation experiments to validate the effectiveness of using Zipformer as the backbone of the flow-matching decoder. As shown in Table V, convolution modules are important in assuring high intelligibility. The U-Net-like downsampling and bypass are consistently helpful in all metrics, showing the effectiveness of the U-Net-like inductive bias on the flow-matching-based TTS model. Notably, removing downsampling but keeping the bypass leads to an architecture similar to the flat Unet-style linked Transformer architecture in [3], which has better results than the model without both downsampling and bypass. And removing bypass connections while retaining downsampling leads to severely degraded results, highlighting the importance of cross-resolution bypass connections in U-net style downsampling structure. Finally, we show the impact of re-using the attention weight. By re-using the attention weight in the NLA module, the performance is consistently improved in all metrics. And the design of sharing attention weight in two self-attention module leads to similar results with the baseline but has improved efficiency.
@@ VI Conclusion
In this paper, we propose ZipVoice, a fast and high-quality flow-matching-based zero-shot TTS model. Leveraging a Zipformer-based backbone for the flow-matching decoder, ZipVoice achieves parameter efficiency while maintaining strong modeling capabilities. An average upsampling strategy, combined with a Zipformer text encoder, ensures stable speech-text alignment and intelligibility. In addition, a flow distillation method is proposed to reduce the NFEs, significantly accelerating inference. Experimental results show that ZipVoice matches SOTA zero-shot TTS models in speech quality but with fewer parameters and faster inference speed.

### v3

@@ Abstract
Existing large-scale zero-shot text-to-speech (TTS) models deliver high speech quality but suffer from slow inference speeds due to massive parameters. To address this issue, this paper introduces ZipVoice, a high-quality flow-matching-based zero-shot TTS model with a compact model size and fast inference speed. Key designs include: 1) a Zipformer-based vector field estimator to maintain adequate modeling capabilities under constrained size; 2) Average upsampling-based initial speech-text alignment and Zipformer-based text encoder to improve speech intelligibility; 3) A flow distillation method to reduce sampling steps and eliminate the inference overhead associated with classifier-free guidance. Experiments on 100k hours multilingual datasets show that ZipVoice matches state-of-the-art models in speech quality, while being 3 times smaller and up to 30 times faster than a DiT-based flow-matching baseline. Codes, model checkpoints and demo samples are publicly available. [^1] https://github.com/k2-fsa/ZipVoice
@@ I Introduction
Recently, flow-matching [12], a variant of diffusion models, has been widely adopted in TTS [2, 3, 4, 13, 14] due to its reduced sampling steps compared to previous diffusion-based models. A representative flow-matching-based zero-shot TTS system is E2-TTS [3]. E2-TTS addresses text-speech alignment by padding text tokens with filler tokens to match the speech length, and letting the Transformer [15]-based vector field estimator to learn the alignment implicitly. This approach significantly simplifies the system architecture and training process by obviating the need for token-level duration prediction. Nevertheless, achieving satisfactory performance with E2-TTS still necessitates a large model size and dozens of sampling steps, which inevitably results in slow inference.
To address the aforementioned challenges, we introduce ZipVoice, a high-quality zero-shot TTS model with a compact model size and fast inference speed. Built upon the flow-matching framework for its simplicity and superior performance, ZipVoice ensure efficient and high-quality speech generation with following designs. Firstly, Zipformer [16] is incorporated as the backbone of vector field estimator to ensure adequate model capacity within a limited parameter budget. We demonstrate that the Zipformer architecture, although originally designed for automatic speech recognition (ASR), is also well-suited for TTS tasks. Secondly, since the practice of padding text tokens with filler tokens in E2-TTS results in suboptimal speech-text alignment and compromised intelligibility [4], to ensure the intelligibility of ZipVoice, on the one hand, we propose an average upsampling approach that assumes uniform token durations within a sentence. This straightforward strategy preserves system simplicity by eliminating the need for token-level duration prediction, while maintaining stable alignment between speech and text. On the other hand, a Zipformer-based text encoder is employed to extract better text token representations. Third, flow-matching-based TTS models typically require many sampling steps to generate high-quality speech, and the widely used classifier-free guidance (CFG) [17] strategy further introduces an additional inference pass without condition in each sampling step. To mitigate this, we design a flow distillation method to distill the pre-trained flow-matching model, thereby reducing the number of sampling steps and avoiding the extra computational overhead introduced by CFG.
Figure 1: Illustration of ZipVoice training (left) and inference (right).
@@ II-A Preliminary: Flow-Matching
After training with the CFM loss, we can generate samples by solving the ordinary differential equation (ODE). Using the Euler solver as an example, starting from an initial sample $x_{0}\sim p_{0}$ (e.g., Gaussian noise), we can iteratively updates the sample towards the target distribution $p_{1}$. For a discrete time sequence $0=t_{0}<\dots<t_{k}<\dots<t_{K}=1$, the Euler solver updates the sample at each step $k$ as:
@@ II-B Overview
The architecture of the ZipVoice model is illustrated in Fig. 1. It consists of two components: a text encoder and a vector field estimator, both employing Zipformer as the backbone. We explain the advantage of this design in subsection II-C.
@@
To enable zero-shot TTS capabilities, we adopt the speech infilling task following [2]. Specifically, a binary temporal mask $m\in\{0,1\}^{D\times T}$, where 1 denotes masked positions, is applied to the speech feature $x_{1}\in\mathbb{R}^{D\times T}$. The model is trained to reconstruct $m\odot x_{1}$ given the speech condition $(1-m)\odot x_{1}$, the interpolated noisy speech feature $x_{t}=(1-t)x_{0}+tx_{1}$, and the text condition $z$. As shown in Fig. 1, these three inputs have the same length and are concatenated along the feature dimension to form the input to the vector field estimator.
@@ II-C ZipVoice Architecture with Zipformer Backbone
ZipVoice comprises a Zipformer-based text encoder and a Zipformer-based vector field estimator, with the latter housing the majority of the model’s parameters. Zipformer is an encoder structure designed for automatic speech recognition (ASR). It distinguishes itself from the standard Transformer structures through three key designs: U-Net-like architecture, convolution modules, and attention weight reuse mechanism.
Zipformer is well-suited as the vector field estimator backbone for the following reasons. Firstly, the U-Net architecture, which processes feature representations at varying resolutions, is widely recognized as an effective inductive bias in diffusion models [18, 19]. Secondly, convolutional neural networks (CNNs) are proficient in capturing fine-grained local feature patterns, complementing the Transformer’s strength in modeling long-range global dependencies. In TTS tasks, the nature of speech leads to strong correlations between adjacent hidden states [20, 21], making CNNs particularly well-suited for enhancing modeling capacity. Finally, Zipformer improves parameter efficiency by reusing attention weights across three modules within a single layer: two self-attention modules and a non-linear attention (NLA) module. This design not only reduces parameter count by sharing projection layers of query and key, but also enhances computational efficiency via attention weight sharing. Therefore, while originally engineered for ASR, these attributes endow Zipformer with ideal compatibility as the backbone of vector field estimators.
Although some recent flow-matching models [3, 4] omit dedicated text encoders by feeding text embeddings directly into the vector field estimator, we find that a well-designed text encoder can improve speech intelligibility. Our Zipformer-based text encoder retains convolutional modules and attention weight reuse while omitting the U-Net structure, aligning with hybrid CNN-Transformer designs in prior works [22, 23].
@@ II-D Speech-Text Alignment with Average Upsampling
Despite its simplicity, this approach suffers from inaccurate alignment and slow convergence  [3]. F5-TTS addresses this by incorporating a ConvNeXt [26] module to refine the padded text condition. ZipVoice instead applies a parameter-free average upsampling strategy to address this issue, under the intuitive assumption of uniform token durations, i.e., each token has identical duration within a sentence. Specifically, when there are $N$ text tokens and $T$ frames speech features frames, the duration of each text token is computed with:
@@
where the floor operation $\lfloor\cdot\rfloor$ ensures the expanded text features do not exceed the speech feature length. Under the practically valid assumption $T\geq N$, the minimum token duration is 1.
Each text embedding is repeated $d$ times and the length of the text feature is expanded from $N$ to $d\cdot N$. If $T>d\cdot N$, the expanded text features is further padded with $T-d\cdot N$ filler embeddings. The final text feature $z\in\mathbb{R}^{F\times T}$ is used as the text condition of the vector field estimator.
@@ II-E Speedup ZipVoice with Flow Distillation
Given a pre-trained TTS model $\theta^{T}$ as the teacher, we initialize the student model $\theta^{S}$ with the parameters of $\theta^{T}$. Since we want the student model to benefit from the CFG inference while avoiding the additional model evaluation of CFG, we modify the student model to be conditioned on the CFG strength [27]. Specifically, the CFG strength $\omega$ is first processed through a Fourier embedding and a linear layer, and then integrated into the model, which is similar to how the time-step is incorporated into the vector field estimator.
@@
The two step sizes, $t_{\text{mid}}-t$ and $t_{\text{dest}}-t_{\text{mid}}$ are uniformly sampled from $0,\Delta t_{\text{max}}]$. $\omega$ is also uniformly sampled between a given interval $\omega_{\text{min}},\omega_{\text{max}}]$. We empirically find that dynamically sampling these values yields better performance than using fixed values.
Then, the teacher vector field is computed as:
@@ II-F Inference Strategy
We use a time-dependent CFG strategy to achieve a better trade-off between speaker similarity and intelligibility: in early NFEs, only the text condition is dropped for unconditional predictions, while in later steps, both text and audio conditions are dropped.
TABLE I: Evaluation results of zero-shot TTS models trained on large-scale datasets. The boldface denotes the best result. * Results from papers. $\S$ Unofficial implementations. $\dagger$ Inference with official checkpoints. $\ddagger$ trained by us with official codes.
@@ IV-B Model
ZipVoice consists of a Zipformer-based text encoder and a Zipformer-based vector field estimator. The text encoder consists of 4 Zipformer layers, each having an encoder dimension of 192, a feedforward dimension of 512. For the vector field estimator, it contains 5 stacks of Zipformer encoder layers operating at $1\times,2\times,4\times,2\times,1\times]$ downsampling rates and with $[2,2,4,4,4]$ layers, respectively. Each layer in the vector field estimator has a embedding dimension of 512 and a feedforward dimension of 1536. The total number of parameters of ZipVoice models are around 123M.
@@ V Experimental Results
V-A Comparison with Baseline Models on Large-Scale Datasets
@@
V-B Comparison with Baseline Models on Small-Scale Datasets
TABLE III: Evaluation results of zero-shot TTS models trained on small-scale datasets. The boldface denotes the best result. $\S$ Unofficial implementations. $\dagger$ Inference with official checkpoints. $\ddagger$ trained by us with official codes.
@@ V-C Ablation Experiments on the ZipVoice Structure
TABLE IV: Ablation studies on the ZipVoice architecture.
@@ V-D Ablation Experiments on the Zipformer Backbone
We conduct detailed ablation experiments to validate the effectiveness of using Zipformer as the backbone of the vector field estimator. As shown in Table V, convolution modules are important in assuring high intelligibility. The U-Net-like downsampling and bypass are consistently helpful in all metrics, showing the effectiveness of the U-Net-like inductive bias on the flow-matching-based TTS model. Notably, removing downsampling but keeping the bypass leads to an architecture similar to the flat Unet-style linked Transformer architecture in [3], which has better results than the model without both downsampling and bypass. And removing bypass connections while retaining downsampling leads to severely degraded results, highlighting the importance of cross-resolution bypass connections in U-net style downsampling structure. Finally, we show the impact of re-using the attention weight. By re-using the attention weight in the NLA module, the performance is consistently improved in all metrics. And the design of sharing attention weight in two self-attention module leads to similar results with the baseline but has improved efficiency.
@@ VI Conclusion
In this paper, we propose ZipVoice, a fast and high-quality flow-matching-based zero-shot TTS model. Leveraging a Zipformer-based backbone for the vector field estimator, ZipVoice achieves parameter efficiency while maintaining strong modeling capabilities. An average upsampling strategy, combined with a Zipformer text encoder, ensures stable speech-text alignment and intelligibility. In addition, a flow distillation method is proposed to reduce the NFEs, significantly accelerating inference. Experimental results show that ZipVoice matches SOTA zero-shot TTS models in speech quality but with fewer parameters and faster inference speed.